OpenStudy (anonymous):

a right triangle with sides ,measuring 6cm , 8cm , and 10cm has a square constructed on each of its three sides. Compare the area of the square on the longest side to the sum of the areas of the two squares on the two shorter legs

OpenStudy (anonymous):

area of the largest square=10^2=100 sum of the areas of the smaller squares= 6^2+8^2=100 so they are equal